OpenStudy (anonymous):

root of 2 in the power of log with the base of 2 of 25. I need to evaluate without calculator. thank you

OpenStudy (anonymous):

\[\log_225^{\sqrt{2}}\rightarrow log_2 5^{(2(\sqrt{2})}\rightarrow 2\sqrt{2}*\log_2(2*(5/4)))\]